Find the product. 
(5 – 2x) (3 + x)
Find the product. 
(5 – 2x) (3 + x)

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

(5 – 2x) (3 + x)
= 5 (3 + x) - 2x (3 + x)
= 15 + 5x - 6x - 2x^2
= 15 - x -2 x^2

Thanks